# Operators’ cognitive performance under extreme hot-humid exposure and its physiological-psychological mechanism based on ECG, fNIRS, and Eye Tracking

### v1

@@ 2   Methodology
performance
Situation awareness
workload
Work memory
task completion rate [%]
SAGAT completion time [s]
mental demand score
number of correct responses [n.u.]
error numbers [n.u.]
total score
physical demand score
number of wrong responses [n.u.]
fault detection time [s]
level1 perception sore
temporal demand score
number of missed responses [n.u.]
fault confirmation time [s]
level2 comprehension score
performance score
mean reaction time of all responses [s]
completion time [s]
level3 projection score
effort score
standard deviation of all response reaction times [s]
communication score
frustration score
mean reaction time of all correct response [s]
total score
standard deviation of all correct response reaction times [s]
@@
Domain
ECG features
Unit
Description
Time domain
RR interval
[s]
Average of time distance between two adjacent R peaks
HR
[bpm]
Numbers of heart beat every minute
R amplitude
No unit
R wave peak amplitude
P amplitude
No unit
P wave peak amplitude
QRS width
[s]
Time between onset and end of the QRS complex
PRQ width
[s]
Time between onset of the P wave to the Q wave
QT interval
[s]
Time between the beginning of the Q wave and the end of the T wave
QTC interval
[s]
QT time interval divided by the square root of the RR interval
ST interval
[s]
Time between the S wave to the end of the T wave
SDNN
[s]
Standard deviation of RR time series
RMSSD
[s]
Root mean square of the difference of all subsequent RR intervals
SDSD
[s]
Standard deviation of the difference of all subsequent RR intervals
PNN50
Percentage of RR intervals in which the change of successive NN exceeds 50 ms
Frequency domain
VLF
[$\mathrm{s}^{2}$]
Spectral power of the RR time series in the band [0; 0.04 Hz]
LF
[$\mathrm{s}^{2}$]
Spectral power of the RR time series in the band [0.04; 0.15 Hz]
HF
[$\mathrm{s}^{2}$]
Spectral power of the RR time series in the band [0.15; 0.4 Hz]
VHF
[$\mathrm{s}^{2}$]
Spectral power of the RR time series in the band [0.4; 3 Hz]
Sympathetic-vagal balance
No unit
Ratio between LF and HF
Sympathetic
No unit
Ratio between LF and (VLF+LF+HF)
Vagal
No unit
Ratio between HF and (VLF+LF+HF)
@@
Domain
Feature
Unit
Description
AOI related
Number of glances$>$2s
No unit
The number of single glance more than 2 seconds
Total Glance time
[s]
Total glance time on a specific AOI
Mean Glance Duration
[s]
Average duration of each glance, total glance time / number of glances
Glance Rate
[1/s]
The number of glances on a AOI every second
AOI Attention Ratio
Glance time on a specific AOI / task duration
Vertical Eye Activity
[pixel]
Standard deviation of y coordinates in the screen
Mean fixation duration left
[ms]
Average fixation duration of left eye in an AOI
Mean fixation duration right
[ms]
Average fixation duration of right eye in an AOI
Number of saccades right
No unit
Total number of saccades of right eye in an AOI
Task and pupil related
Duration
[s]
Duration on an interest page
Percentage Transition Times
Time of transiting between different AOIs / duration
Pupil left avg
Average pupil diameter of left eye
Pupil right avg
Average pupil diameter of right eye
@@ 3   Results
error counts
2
4.00±6.91
3.41±5.65
4.78±6.95
2.89±2.47
0.21
0.02
0.10
fault detection time
1
3.71±2.20
2.82±1.13
3.03±1.41
3.14±2.12
0.05
0.37
0.27
fault confirm time
1
33.33±14.35
28.65±12.65
26.67±10.99
25.35±8.68
$<$0.01
0.18
0.91
completion time
82.76±37.41
73.64±31.12
70.91±23.18
61.54±18.38
0.05
0.37
$<$0.01
@@
Total score
1
82.59±21.14
81.85±24.07
86.30±18.94
80.37±26.31
0.82
0.18
0.03
comprehension
2
23.33±4.16
23.89±3.49
22.96±3.74
23.70±2.97
0.08
0.02
0.16
communication
1
4.44±1.60
4.07±1.98
4.63±1.33
3.89±2.12
0.33
0.08
0.04
@@
physical demand
23.68±19.11
24.18±17.50
30.17±20.14
43.56±22.81
0.56
0.01
$<$0.01
temporal demand
43.74±22.57
38.26±19.44
38.37±22.16
41.73±22.06
0.03
0.97
0.16
effort
44.63±23.86
39.08±22.10
40.93±22.00
46.94±22.32
0.09
0.33
0.03
total score
48.40±29.30
42.78±22.55
43.42±23.50
46.47±22.35
0.03
0.72
0.09
@@
missed number
0.62±0.90
0.27±0.53
0.73±0.96
0.92±2.56
0.01 0.02
0.73
mean reaction time of correct responses
1.12±0.18
1.10±0.16
1.07±0.17
1.02±0.16
0.39
0.19
$<$0.01
@@
RR interval [s]
0.68±0.08
0.69±0.08
0.67±0.08
0.61±0.08
1.00
0.34
$<$0.01
0.27
$<$0.01
$<$0.01
HR [bpm]
89.29±11.08
89.05±10.91
91.42±11.32
100.94±13.77
1.00
0.45
$<$0.01
0.35
$<$0.01
$<$0.01
QT interval [s]
0.41±0.13
0.41±0.13
0.40±0.13
0.36±0.13
1.00
0.99
0.03
0.96
0.02
0.06
ST interva [s]
0.33±0.13
0.34±0.13
0.33±0.13
0.29±0.13
1.00
0.99
0.02
0.96
0.01 0.04
SDNN [s]
0.04±0.01
0.04±0.01
0.03±0.01
0.03±0.02
0.78
0.25
$<$0.01
0.80
0.01
0.14
RMSSD [s]
0.02±0.01
0.02±0.01
0.02±0.01
0.02±0.02
0.77
0.01
$<$0.01
0.16
$<$0.01
0.43
SDSD [s]
0.02±0.01
0.02±0.01
0.02±0.01
0.02±0.02
0.77
0.01
$<$0.01
0.16
$<$0.01
0.43
PNN50 [%]
2.26±3.08
2.04±2.97
1.38±2.12
0.92±1.47
0.88
0.02
$<$0.01
0.14
$<$0.01
0.44
@@
Sympathetic-vagal balance
1.84±1.41
2.38±2.08
2.79±2.36
2.89±3.30
0.24
0.01
$<$0.01
0.50
0.29
0.98
Vagal
0.38±0.20
0.34±0.19
0.30±0.18
0.29±0.20
0.28
$<$0.01
$<$0.01
0.29
0.15
0.99
@@
Feature
Task
C1
C2
C3
C4
p C2-C1
p C3-C2
p C4-C3
Number of glances$>$2s
1
0.33±0.79
0.19±0.47
0.67±1.47
0.65±1.30
0.36
0.04
0.76
Total Glance time [s]
3
1.07±0.40
1.01±0.39
1.30±0.52
1.02±0.38
0.46
0.04 0.02
Mean Glance Duration [s]
3
0.95±0.39
0.86±0.38
1.21±0.53
0.96±0.43
0.46
0.01 0.04
Glance Rate [1/s]
3
0.05±0.02
0.06±0.02
0.05±0.01
0.06±0.03
0.02
0.09
0.28
AOI Attention Ratio [%]
1
60.50±15.31
61.10±15.95
58.75±18.20
51.87±22.15
0.82
0.31
$<$0.01
AOI Attention Ratio [%]
3
4.79±1.66
4.92±1.57
6.32±2.26
5.36±1.71
0.72
0.02
0.08
Vertical Eye Activity [pixel]
1
80.66±22.55
88.18±32.22
82.66±33.15
99.87±51.64
0.16
0.18
0.01
Mean fixation duration left [ms]
2
1172.75±565.56
1592.61±1014.11
1702.61±1314.78
1404.03±929.85
0.05
0.72
0.27
Mean fixation duration right [ms]
1
439.73±97.46
472.45±131.91
475.96±96.15
432.08±121.18
0.04
0.88
0.03
Mean fixation duration right [ms]
3
857.67±484.26
781.05±430.88
1039.89±676.90
664.45±392.81
0.99
0.08
0.04
Mean saccade angle right [deg]
1
3.88±1.97
4.70±0.94
4.90±0.95
4.99±1.38
0.02
0.32
0.79
Number of saccades right
1
6.93±5.17
6.94±3.60
8.03±5.40
10.62±8.37
0.67
0.30
0.05
@@
Feature
Task
C1
C2
C3
C4
p C2-C1
p C3-C2
p C4-C3
Duration [s]
3
22.62±4.56
20.61±4.73
20.67±4.06
19.03±2.74
0.02
0.94
0.04
Percentage Transition Times [%]
1
31.12±12.07
30.25±11.89
30.36±11.43
36.10±16.72
0.71
0.95
0.02
Pupil left avg
2
765.62±227.20
767.62±214.42
727.42±200.13
723.69±218.04
0.86
$<\textbf{0.01}$
0.98
Pupil right avg
1
804.64±154.96
785.83±179.94
770.75±199.82
816.48±258.55
0.01
0.27
0.01
Pupil right avg
2
806.17±145.28
802.25±179.40
782.20±185.46
816.61±252.90
0.02
0.14
0.05
Pupil right avg
3
782.45±147.29
768.69±183.24
762.66±209.55
788.31±275.54
0.02
0.70
0.26

### v2

@@ 2   Methodology
performance Situation awareness workload Work memory
task completion rate [%] SAGAT completion time [s] mental demand score number of correct responses [n.u.]
error numbers [n.u.] total score physical demand score number of wrong responses [n.u.]
fault detection time [s] level1 perception sore temporal demand score number of missed responses [n.u.]
fault confirmation time [s] level2 comprehension score performance score mean reaction time of all responses [s]
completion time [s] level3 projection score effort score standard deviation of all response reaction times [s]
communication score frustration score mean reaction time of all correct response [s]
total score standard deviation of all correct response reaction times [s]
@@
Domain ECG features Unit Description
Time domain RR interval [s] Average of time distance between two adjacent R peaks
HR [bpm] Numbers of heart beat every minute
R amplitude No unit R wave peak amplitude
P amplitude No unit P wave peak amplitude
QRS width [s] Time between onset and end of the QRS complex
PRQ width [s] Time between onset of the P wave to the Q wave
QT interval [s] Time between the beginning of the Q wave and the end of the T wave
QTC interval [s] QT time interval divided by the square root of the RR interval
ST interval [s] Time between the S wave to the end of the T wave
SDNN [s] Standard deviation of RR time series
RMSSD [s] Root mean square of the difference of all subsequent RR intervals
SDSD [s] Standard deviation of the difference of all subsequent RR intervals
PNN50 % Percentage of RR intervals in which the change of successive NN exceeds 50 ms
Frequency domain VLF [$\mathrm{s}^{2}$] Spectral power of the RR time series in the band [0; 0.04 Hz]
LF [$\mathrm{s}^{2}$] Spectral power of the RR time series in the band [0.04; 0.15 Hz]
HF [$\mathrm{s}^{2}$] Spectral power of the RR time series in the band [0.15; 0.4 Hz]
VHF [$\mathrm{s}^{2}$] Spectral power of the RR time series in the band [0.4; 3 Hz]
Sympathetic-vagal balance No unit Ratio between LF and HF
Sympathetic No unit Ratio between LF and (VLF+LF+HF)
Vagal No unit Ratio between HF and (VLF+LF+HF)
@@
Domain Feature Unit Description
AOI related Number of glances$>$2s No unit The number of single glance more than 2 seconds
Total Glance time [s] Total glance time on a specific AOI
Mean Glance Duration [s] Average duration of each glance, total glance time / number of glances
Glance Rate [1/s] The number of glances on a AOI every second
AOI Attention Ratio [%] Glance time on a specific AOI / task duration
Vertical Eye Activity [pixel] Standard deviation of y coordinates in the screen
Mean fixation duration left [ms] Average fixation duration of left eye in an AOI
Mean fixation duration right [ms] Average fixation duration of right eye in an AOI
Number of saccades right No unit Total number of saccades of right eye in an AOI
Task and pupil related Duration [s] Duration on an interest page
Percentage Transition Times [%] Time of transiting between different AOIs / duration
Pupil left avg Average pupil diameter of left eye
Pupil right avg Average pupil diameter of right eye
@@ 3   Results
error counts 2 4.00±6.91 3.41±5.65 4.78±6.95 2.89±2.47 0.21 0.02 0.10
fault detection time 1 3.71±2.20 2.82±1.13 3.03±1.41 3.14±2.12 0.05 0.37 0.27
fault confirm time 1 33.33±14.35 28.65±12.65 26.67±10.99 25.35±8.68 $<$0.01 0.18 0.91
completion time / 82.76±37.41 73.64±31.12 70.91±23.18 61.54±18.38 0.05 0.37 $<$0.01
@@
Total score 1 82.59±21.14 81.85±24.07 86.30±18.94 80.37±26.31 0.82 0.18 0.03
comprehension 2 23.33±4.16 23.89±3.49 22.96±3.74 23.70±2.97 0.08 0.02 0.16
communication 1 4.44±1.60 4.07±1.98 4.63±1.33 3.89±2.12 0.33 0.08 0.04
@@
physical demand / 23.68±19.11 24.18±17.50 30.17±20.14 43.56±22.81 0.56 0.01 $<$0.01
temporal demand / 43.74±22.57 38.26±19.44 38.37±22.16 41.73±22.06 0.03 0.97 0.16
effort / 44.63±23.86 39.08±22.10 40.93±22.00 46.94±22.32 0.09 0.33 0.03
total score / 48.40±29.30 42.78±22.55 43.42±23.50 46.47±22.35 0.03 0.72 0.09
@@
missed number / 0.62±0.90 0.27±0.53 0.73±0.96 0.92±2.56 0.01 0.02 0.73
mean reaction time of correct responses / 1.12±0.18 1.10±0.16 1.07±0.17 1.02±0.16 0.39 0.19 $<$0.01
@@
RR interval [s] 0.68±0.08 0.69±0.08 0.67±0.08 0.61±0.08 1.00 0.34 $<$0.01 0.27 $<$0.01 $<$0.01
HR [bpm] 89.29±11.08 89.05±10.91 91.42±11.32 100.94±13.77 1.00 0.45 $<$0.01 0.35 $<$0.01 $<$0.01
QT interval [s] 0.41±0.13 0.41±0.13 0.40±0.13 0.36±0.13 1.00 0.99 0.03 0.96 0.02 0.06
ST interva [s] 0.33±0.13 0.34±0.13 0.33±0.13 0.29±0.13 1.00 0.99 0.02 0.96 0.01 0.04
SDNN [s] 0.04±0.01 0.04±0.01 0.03±0.01 0.03±0.02 0.78 0.25 $<$0.01 0.80 0.01 0.14
RMSSD [s] 0.02±0.01 0.02±0.01 0.02±0.01 0.02±0.02 0.77 0.01 $<$0.01 0.16 $<$0.01 0.43
SDSD [s] 0.02±0.01 0.02±0.01 0.02±0.01 0.02±0.02 0.77 0.01 $<$0.01 0.16 $<$0.01 0.43
PNN50 [%] 2.26±3.08 2.04±2.97 1.38±2.12 0.92±1.47 0.88 0.02 $<$0.01 0.14 $<$0.01 0.44
@@
Sympathetic-vagal balance 1.84±1.41 2.38±2.08 2.79±2.36 2.89±3.30 0.24 0.01 $<$0.01 0.50 0.29 0.98
Vagal 0.38±0.20 0.34±0.19 0.30±0.18 0.29±0.20 0.28 $<$0.01 $<$0.01 0.29 0.15 0.99
@@
Feature Task C1 C2 C3 C4 p C2-C1 p C3-C2 p C4-C3
Number of glances$>$2s 1 0.33±0.79 0.19±0.47 0.67±1.47 0.65±1.30 0.36 0.04 0.76
Total Glance time [s] 3 1.07±0.40 1.01±0.39 1.30±0.52 1.02±0.38 0.46 0.04 0.02
Mean Glance Duration [s] 3 0.95±0.39 0.86±0.38 1.21±0.53 0.96±0.43 0.46 0.01 0.04
Glance Rate [1/s] 3 0.05±0.02 0.06±0.02 0.05±0.01 0.06±0.03 0.02 0.09 0.28
AOI Attention Ratio [%] 1 60.50±15.31 61.10±15.95 58.75±18.20 51.87±22.15 0.82 0.31 $<$0.01
AOI Attention Ratio [%] 3 4.79±1.66 4.92±1.57 6.32±2.26 5.36±1.71 0.72 0.02 0.08
Vertical Eye Activity [pixel] 1 80.66±22.55 88.18±32.22 82.66±33.15 99.87±51.64 0.16 0.18 0.01
Mean fixation duration left [ms] 2 1172.75±565.56 1592.61±1014.11 1702.61±1314.78 1404.03±929.85 0.05 0.72 0.27
Mean fixation duration right [ms] 1 439.73±97.46 472.45±131.91 475.96±96.15 432.08±121.18 0.04 0.88 0.03
Mean fixation duration right [ms] 3 857.67±484.26 781.05±430.88 1039.89±676.90 664.45±392.81 0.99 0.08 0.04
Mean saccade angle right [deg] 1 3.88±1.97 4.70±0.94 4.90±0.95 4.99±1.38 0.02 0.32 0.79
Number of saccades right 1 6.93±5.17 6.94±3.60 8.03±5.40 10.62±8.37 0.67 0.30 0.05
@@
Feature Task C1 C2 C3 C4 p C2-C1 p C3-C2 p C4-C3
Duration [s] 3 22.62±4.56 20.61±4.73 20.67±4.06 19.03±2.74 0.02 0.94 0.04
Percentage Transition Times [%] 1 31.12±12.07 30.25±11.89 30.36±11.43 36.10±16.72 0.71 0.95 0.02
Pupil left avg 2 765.62±227.20 767.62±214.42 727.42±200.13 723.69±218.04 0.86 $<\textbf{0.01}$ 0.98
Pupil right avg 1 804.64±154.96 785.83±179.94 770.75±199.82 816.48±258.55 0.01 0.27 0.01
Pupil right avg 2 806.17±145.28 802.25±179.40 782.20±185.46 816.61±252.90 0.02 0.14 0.05
Pupil right avg 3 782.45±147.29 768.69±183.24 762.66±209.55 788.31±275.54 0.02 0.70 0.26